OpenStudy (keyonsay):

The area of the circle above is 1,519.76 units2. What is the circumference of the circle? (Use 3.14 for .)

OpenStudy (anonymous):

\[A=\pi r^{2}=3.14r^{2}\] A is given as 1519.76, so \[1519.76=3.14r^{2}\] Isolating r: \[r^{2}=1519.76/3.14=484\] Taking the square root of both sides gives \[r=22\] The circumference of a circle, given its radius, is \[C=2\pi r=2*3.14*22=138.16\] So, the circumference of the circle is 138.16 units.

OpenStudy (anonymous):

\[A= \pi r ^{2}\] A=Area of a circle r=radius \[C = 2 \pi r\] C=Circumference r=radius

OpenStudy (keyonsay):

thank you